Add without using number line:
(а) 11 + (-7)
(b) (-13) + (+18)
(c) (-10) + (+19)
(d) (-250) + (+150)
(e) (-380) + (-270)
(f) (-217) + (-100).

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Solution: (a) 11 + (-7) = 4 + (+7) + (-7) [∵ (+7) + (-7) = 0] = 4 + 0 = 4 Hence, 11 + (-7) = 4. (b) (-13) + (+18) = (-13) + (+13) + (+5) [∵ (-13) + (+13) = 0] = 0 + (+ 5) = 5 Hence, (- 13) + (+ 18) = 5. (c) (-10) + (+19) = (-10) + (+10) + (+9) [∵ (-10) + (10) = 0] = 0 + (+9) = 9 Hence, (-10) + (19) = 9. (d) (-250) + (+150) = (-100) + (-150) + (+150) = (-100) + 0 = -100 [∵(-150) + (+150) = 0] Hence, (-250) + (+150) = – 100. (e) (-380) + (-270) = – [380 + 270] = (-650) Hence, (-380) + (-270) = (-650). (f) (-217) + (-100) = – [217 + 100] = -317